OpenStudy (doughnuttable):

If I have the solution u(x,t)=cos(t) sin(x), how would I show that this solution also has the form u(x,t) = p(x-ct) + q(x+ct)?

OpenStudy (doughnuttable):

It looks like you can reduce the order of the equation using \( \left\{ \begin{array}{11} u_t + cu_x = w & 1)\\ \ w_t - c w_x = 0 & 2) \end{array} \right.\) When you solve this, you get one part of the solution.: \( \begin{align*} \frac{dt}{dx} &= \frac{1}{-c} \\ -cdt &= dx \\ 0 &= cdt + dx \\ k &= ct + x \\ f(k) &= f(ct +x), & f(k) \mbox{ is a solution to 2) } \\ u_t + cu_x &= \color{green} {w = f(ct+x) }\\ f(ct+x) &= -\sin t \sin x + c \cos t \cos x \end{align*} \)

OpenStudy (doughnuttable):

Then you do the same for \( \left\{ \begin{array}{11} u_t - cu_x = v & 1)\\ \ v_t + c v_x = 0 & 2) \end{array} \right. \) And add the two parts together. I get \( u(x,t) = -\sin t \sin x + c \cos t \cos x -\sin t \sin x -c \cos t \cos x\)

OpenStudy (doughnuttable):

Whoops, looks like that's wrong.

OpenStudy (doughnuttable):

You need to isolate \(u_t\) and \(u_x\) by adding and subtracting \(f(x+ct)\) and \( g(x−ct) \). Then integrate, then add the results together.